A Doppler flow meter uses ultrasound waves to measure blood-flow speeds. Suppose the device emits sound at 3.9MHz , and the speed of sound in human tissue is about 1540 m/s. What is the expected beat frequency if blood is flowing in large leg arteries at 3.0

cm/s directly away from the sound source?
Answer #1

f = f0*(c-vr)/(c+vr) = 3900*(300000000-1540*0.03)/(300000000+1540*03) =3899.998 Hz
Beat frequency df = f0-f = 3900Hz-3899.998Hz=0.040039 Hz